A particle starts from rest and accelerates uniformly in a straight line. The particle travels a distance of 6.70 m in 2.50 s .

How fast, in miles/hour (mph), will the particle be traveling after a time of 4.50 s has elapsed? Assume 1 mile=1609.34 m.

Express your answer using three significant figures.
Answer #1

Since particle starts from rest ,initial velocity is zero

Distance traveled in 2.5 s = 6.7m

Use formula sut +1/2at2 to find acceleration

6.7m (Om/s) (2.5s) 1/2a(2.5s)

6.702.50.5 a 2.52

6.7 = 3.125a

a = 2.144m

-------------------------

Use formula v=u+at to find speed after 4.5s

u = Orn/s+ (2.144m/ s-) (4.5s)

u = 0 + 2.144 * 4.5

9.648m/s

--------------------

Given 1 mile=1609.34 m

1 hr =3600s

9.648m/s -9.648m/s*1 *1

m 1mile3600s s 1609.34m 1hr 9.648m/s-9.648-

mile 9.648m/s 9.648 * 3600 * 1609.34 r

9.648m/s=21.582mi/hr

ANSWER: 21.6mi/hr

===================